OpenStudy (anonymous):

Which of the following is equal to square root of the cube root of 2? 2 to the power of 2 over 3 2 to the power of 3 over 2 2 to the power of 1 over 3 2 to the power of 1 over 6

OpenStudy (anonymous):

@aaronq @Kainui @PaulaLovesSchool13 @confluxepic @sweetburger @iambatman

OpenStudy (anonymous):

I'll see if I could find a link to help you. The others can explain it to you :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

It's a link to help you :)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

@logancoop @mathstudent55 @mathmath333 @LUCIFERSKNIGHT @OpiGeode @MrNood @undeadknight26

OpenStudy (anonymous):

Whats goin on

OpenStudy (anonymous):

what u mean

OpenStudy (mrnood):

\[\sqrt{\sqrt[3]{2}}\]

OpenStudy (mrnood):

\[\sqrt{x} = x^{\frac{ 1 }{ 2 }}\] \[\sqrt[3]{x} = x ^{\frac{ 1 }{ 3 }}\] so use properties of exponents to combine those two ....

OpenStudy (mrnood):

\[(x ^{a})^{b} = x ^{ab}\]